3. If light is passed through a double-slit opening onto a screen, the pattern produced on the screen is ___________



A. a bright central band of light with slightly-diminished alternating bright and dark bands called nodes and antinodes.


B. a bright contral band of light with tiny lines toward the edges of the screen


C. a large circle of light with tiny circles around it.


D. equally-sized concentric circles of light



4. Light passes through a double-slit, producing bright and dark hands on a screen The dark hands are caused by waves from each slit _________



A. interfening constructively


B. meeting exactly one wavelength out of phase


C. meeting exactly one half wavelength out of phase.


D. meeting exactly one fourth wavelength out of phase.



5. Light of wavelength 580 nm is incident on a sit of width 0.300 mm: An observing Streen is placed 2.00m from the slit. Find the position of the first order dark frings from the center of the screen.



A. 0.26 mm


B. 1.9 mm


C. 3.9 mm


D. 7.7 mm

1
Expert's answer
2021-12-29T12:34:10-0500

3. If light is passed through a double-slit opening onto a screen, the pattern produced on the screen is a bright central band of light with slightly-diminished alternating bright and dark bands called nodes and antinodes.


4. Light passes through a double-slit, producing bright and dark hands on a screen. The dark hands are caused by waves from each slit meeting exactly one half wavelength out of phase.


5. The position of the first order dark fringes from the center of the screen:


"d \sin θ = mλ,\\\space\\
\theta=\arcsin\dfrac{m\lambda}{d},\\\space\\
\theta=\arcsin\dfrac{(1)(580·10^{-9})}{0.3·10^{-3}}=0.111°."


The position is


"x=L\tan\theta=2\tan0.1°=3.87·10^{-3}\text{ m}."